OpenStudy (anonymous):

Simplify 8(2x^3− 3x + 2) − 12(3x^3 + 2x) + 3(4x − 3)

OpenStudy (anonymous):

i have to sign off if you can please anser the question, thank you.

OpenStudy (anonymous):

16x^3-24x+16-36x^3+24x+12x-9

OpenStudy (anonymous):

rest u can solve

OpenStudy (anonymous):

\[16x^3-24x+16-36x^3+24x+12x-9\]

Parth (parthkohli):

(16x^3 - 24x +16) - (36x^3 + 24x) + (12x - 9) = 16x^3 - 24x + 16 - 36x^3 - 24x + 12 - 9 = -20^3 - 48x + 19

Parth (parthkohli):

\[-20^{3} - 48x + 19\]

OpenStudy (anonymous):

\[= -20^3 - 48x + 19\]